A trader would incur a loss of 15%, if he sells pens at Rs.10.20 per pen. If 12.5% profit is to be earned after allowing 10% discount on the marked price then find the marked price?
Answer by ghostrecog(4)

Discount % =10%
Given loss% 15% and S.P.=10.20
C.P.=SP*100/100-loss%
=10.20*100/100-15
=1020/85
C.P. =12
Given profit % 12.5% and need to find marked price.
When C.P. and profit % are given then S.P.=CP* (100+Profit %) /100
=12*(100+12.5)/100
=12*112.5/100
S.P.=13.5
Let x be the M.P
When S.P. and Discount % are given, M.P.=(S.P.*100)/100-Discount %
x=(13.5*100)/100-10
=1350/90
=15.
The marked price is 15.